# Rhetorical XAI: Explaining AI’s Benefits as well as its Use via Rhetorical Design

### v1

@@ B.3. Reviewed articles (continued)
Kocielnik et al. (2019) Isolated presentation; Expectation management Logical: Subjective metrics User understanding of AI model, Perceived control over system behaviors
Emotional: Subjective metrics Perceived system accuracy based on prior expectations
Liu et al. (2021) Isolated presentation Logical: Objective metrics Appropriate trust; Subjective metrics Real-time assistant useful, Training useful
Ma et al. (2023) Adaptive presentation; Technical assurance Logical: Human-AI agreement; Subjective metrics Perceived complexity, Future use, Perceived usefulness, Perceived helpfulness, Acceptable; Qualitative themes Participants in adaptive workflow were forced to think independently
Credible: Subjective metrics Trust in AI, Confidence, Explanation satisfaction, Trust in Correctness Likelihood; Qualitative themes Some participants distrust and ignore correctness likelihood but most participants referred to or were influenced by correctness likelihood based adaptation.
Miranda et al. (2021) Isolated presentation; Explanation as retriever Logical: Subjective metrics Model behavior understanding, Tool usefulness, Usability.
Nguyen et al. (2018) Adaptive presentation Logical: Qualitative themes Difficult to understand;
Emotional: Qualitative themes Enjoy playing with the tool
B.4. Reviewed articles (continued)
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Okolo et al. (2024) Cultural relatedness; Emotional cues Logical: Qualitative themes Color confused the interpretation of feature importance;
Credible: Qualitative themes Misinterpretation of patient severity due to default feature colors that conflict with medical domain norms;
Emotional: Qualitative themes Negative emotional perception
Okoso et al. (2025) Authority framing; Emotional cues Credible: Subjective metrics Acceptance of AI advice, Qualitative themes Some participants said tome influence their decisions and informal tone was inappropriate for a serious topic;
Emotional: Qualitative themes Some participants expressed discomfort with AI mimicking human emotions
Schemmer et al. (2023) Isolated presentation Logical: Objective metrics Appropriate reliance
Schuff et al. (2022) Isolated presentation Logical: Subjective metrics Self-reported importance ratings of features
Sun et al. (2023a) Isolated presentation; Explanation as retriever Logical: Subjective metrics Model understanding, Tool usability; Qualitative themes Insufficient description about reasonability matrix, Ill-formed local explanation visualization, Attention adjustment can be subjective
Suresh et al. (2022) Isolated presentation; Explanation as retriever Logical: Objective metrics Human-AI agreement; Qualitative themes Nearest neighbors enable reasoning with clinically-relevant concepts and help characterize model uncertainty, Feature-based explanations are difficult to extract high-level, clinically-relevant concepts;
Credible: Qualitative themes Visualizing variation helps assess prediction reliability
van der Waa et al. (2021) Isolated presentation Logical: Objective metrics Advice prediction, Decision factor identification; Subjective metrics Perceived system understanding, Perceived prediction of correctness
Vasconcelos et al. (2023) Isolated presentation Logical: Objective metrics Appropriate reliance, Task difficulty;
Credible (authors’ post-hoc reflections): Stakes of the task;
Emotional (authors’ post-hoc reflections): Enjoyability
Wambsganss et al. (2020) Adaptive & Mechanistic presentation; Explanation as recommender Logical: Subjective metrics Intention to use, Perceived usefulness, Perceived ease of use; Qualitative themes Positive feeling about adaptive learning given its fast and direct feedback, graph-like visualization, and score. However, sometimes the system is less accurate in claim and premise classification.
B.5. Reviewed articles (continued)
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Wang et al. (2019) Mechanistic presentation Logical: Qualitative themes Counterfactual reasoning is useful;
Credible: Qualitative themes Participants felt less trustworthy without the patient context presented in the explanations
Wang and Yin (2021) Isolated presentation Logical: Objective metrics Human understanding of model behaviors, Appropriate trust; Subjective metrics Self-reported understanding of model behaviors
Wang and Yin (2023) Adaptive presentation Logical: Objective metrics Frequency of accepting model decisions; Subjective metrics Perceived accuracy;
@@ B.5. Reviewed articles (continued)
Credible: Subjective metrics Self-reported trust gain, Satisfaction gain
Wang et al. (2023b) Mechanistic presentation; Explanation as recommender; Emotional cues Logical: Subjective metrics Easy to use, Easy to understand, Help understand model, Help find ways to improve plans, Will apply loan again, Usefulness of specific interactions; Qualitative themes Controllable, Requiring small changes or less compromise, Beneficial for life;
Emotional: Subjective metrics Enjoyable to use
Yang et al. (2020) Isolated presentation Logical: Objective metrics Appropriate trust; Subjective metrics Explanation helpfulness, Usability;
@@
Credible: Subjective metrics Self-confidence, Trust meter before and after AI feedback
Zhang et al. (2020) Isolated presentation; Technical assurance Logical: Objective metrics Human-AI agreement

### v2

@@ Appendix B Narrative Review Process and Corresponding Coding Scheme for Reviewed Articles
Logical Self-reported understanding of AI model, perceived control over system behaviors:
p8: “Comparing conditions with and without Explanation revealed a significant positive impact…with higher average level of perceived understanding (mean of both understanding questions) for participants who saw the Explanation than for those who did not…We therefore consider H2.3 supported as the Control Slider intervention significantly increased participants’ perceived level of control over the Scheduling Assistant’s behavior.
Emotional Expected satisfaction of system performance:
p10: Participants in the High Recall version of the system reported no significant difference in satisfaction between the Baseline and Pure-techniques. At the same time, in the High Precision version, participants exposed to Pure-techniques reported significantly higher satisfaction than those in the Baseline…
Liu et al. (2021) How AI Works: Isolated presentation with interactivity
Logical Effects of explanations on appropriate trust (reliance):
p22: “In general, we do not find significant impact from interactive explanations with respect to the performance of human-AI team or human agreement with wrong AI predictions, compared to static explanations. However, humans are more likely to find AI assistance useful with interactive explanations than static explanations in ICPSR and COMPAS, but not in BIOS…
Self-reported real-time assistance useful:
p25: In COMPAS, Interactive/Interactive achieves a significantly higher human perception of real-time assistance usefulness than both Static/Static and Interactive/Static…Perception of Interactive/Static is also significantly higher than that of Static/Static. We find similar results in ICPSR except that the difference between Static/Static and Interactive/Static is not significant…
Self-reported identification of important features:
p25: “These results suggest that participants with interaction are more likely to fixate on demographic features and potentially reinforce human biases,11 but are less likely to identify computationally important features in ICPSR and COMPAS.
Ma et al. (2023) How AI Works: Adaptive presentation; Why AI Merits Use: Technical assurance
Logical Effects of explanation on human-AI agreement (appropriate trust):
Figure 6 shows the human-AI agreement in different human-AI CL situations under three conditions. Results showed that all three conditions made people’s agreement with AI significantly higher when AI’s CL was higher than that when the human’s CL was higher…the human-AI agreement in AI Confidence was significantly higher than Adaptive Workflow , and Adaptive Recommendation, and it was marginally higher than Direct Display. Note that when AI is wrong, a lower agreement with AI is better. These results suggest humans’ less over-trust in AI in our proposed CL exploitation conditions…However, when the AI’s recommendation was correct, we did not observe significant differences in human-AI agreement between the three proposed conditions and AI Confidence baseline.
Qualitative feedback:
Participants in Adaptive Workflow and Adaptive Recommendation were forced to think independently.
Credible Qualitative feedback of correctness likelihood:
Some participants doubted the displayed CL information and ignored it in the decision-making process…Most participants referred to or were influenced by the dis- played CL or CL-based adaptation
Nguyen et al. (2018) How AI Works: Isolated presentation with interactivity
Logical Effect of presenting AI predictions on fact-checking accuracy:
p194: “We observe that there are larger changes for errors made by participants in group Control: those who have not seen our stance predictions change their answers more than those who have…This roughly corresponds to the errors by the veracity classifier, showing again that system predictions can both help users (when correct) or lead users to errors that reflect model fallibility or biases implicit in training data.
Qualitative feedback:
p196: “It was very hard to understand. It seems on one task, I was 100% sure it was true and I was told it was false, I even followed links to verify the sources
Emotional Qualitative feedback:
p196: “…thought it was really cool! I’d enjoy playing with this more if it wasn’t during my work time…
Okolo et al. (2024) How AI Works: Isolated presentation, Cultural relateness, Emotional cues
Logical Qualitative feedback:
p11: “While these interpretations made sense to the CHWs, they were not what the XAI methods intended to convey (they intended to show the original image for reference and the annotated image for explainability, with no storyline among the images). Only a few CHWs could partially understand the explanations with the help of higher-level features such as colors, images, or shapes” p17: “This interpretation was incorrect because the underlying model relied on specific bodily regions (in alignment with the Kramer’s Rule chart in Table 1) to provide a severity prediction, not on the number of yellow boxes.
Credible Qualitative feedback:
p13: “P04 and P05 emphasized that their patients trust them more when they use visual aids. They gave the example of paper-based visual aids they carry to educate pregnant women about when to seek urgent care. When we asked if CHWs would prefer similar paper-based visual aids for jaundice, P04 and P05 preferred the AI-driven app over paper-based visual aids, stating that the app could diagnose the severity of diseases, which the paper-based aids could not.
Emotional Qualitative feedback:
p8: “They were worried that the red color might confuse the participants since it is often used in public health messaging to imply ‘danger’ in the medical sense.” p14: “CHWs often relied on their prior experience with colors in visual aids in the domain of Public Health to interpret what the colors in the XAI visualizations might mean. They had strong color associations, believing that ‘green means safe, yellow means warning, red means danger and gray/blue means lack of blood.’ These color associations largely influenced how they interpreted the visualizations depicting a jaundiced baby.
Okoso et al. (2025) How AI Works: Authority framing, Emotional cues
Credible Self-reported acceptance of AI advice:
In the Formal, Authoritative, and Humorous groups, an upward trend was observed in d tone u, indicating that participants tended to increase their scores. Although no significant overall effect of tone was observed, the tone could possibly affect participants differently depending on specific user attributes.
Qualitative feedback:
Among the 135 participants who noticed a change in tone and provided relevant responses, 59 participants said it did not affect their decisions, whereas 55 participants said it did. Some participants felt that an informal tone was inappropriate for a serious topic, such as recidivism risk, whereas others found the tone easier to understand.
Emotional Qualitative feedback:
A total of 110 participants (37%) expressed either positive or negative opinions about the explanations but ultimately stated that it did not influence their decision. Most felt that the explanations did not provide sufficient information relevant to their needs…those movies based solely on genre and did not consider the explanation, whereas others remarked that it felt unnatural for an AI, which supposedly lacks emotions, to promote a movie
Schemmer et al. (2023) How AI Works: Isolated presentation
Logical Effects of explanation on appropriate reliance:
p418: “In the XAI condition, we can observe a significant increase in relative AI reliance (PAIR) from 29.59% to 38.87% while the (relative self-reliance) RSR does not change significantly (71.87%) for the control condition and 69.45% for the feature importance condition. This means explanations of AI decisions can reduce the share of under-reliance. It is important to highlight that PAIR is not increased simply by relying more often on AI advice, as this would have also reduced the RSR significantly.
Schuff et al. (2022) How AI Works: Isolated presentation
Logical Self-reported importance of features reflect bias of understanding explanations:
p619: “Overall, three studies confirmed the presumably biasing influence of word length, (pairs of) two studies respectively confirmed the effect of sentence length, display index and saliency rank, and one study (each) found significant effects of word position, sentiment polarity, word frequency, capitalization and dependency relation
Sun et al. (2023a) How AI Works: Isolated presentation; Why AI Merits: Explanation as retriever
Logical Improve model performance by using local explanation-based attention steering and evaluate the model vulnerability:
p19-20: “S2 results suggest that (1) the workflow of the local explanation-based attention steering provided a diverse perspective in diagnosing model vulnerability, (2) the direct steering design helped the process of model revision straightforward, and (3) every participant enjoyed improved key model performance measures…The behavioral data we collected shows that all participants generated the model that outperforms (1) its model accuracy, (2) the overlap between the model’s focus and the relevant object types (IoU), and (3) the proportion of reasonable attention out of all images in a test set…After completing the user studies, the majority of users strongly agree that adjusting local explanations can effectively improve model performance.
Qualitative feedback:
p20: “During interviews, all participants shared their positive impressions about the effectiveness of attention adjustment in improving model accuracy, which is the primary objective of conducting model fine-tuning. They also confirmed that the impact of contextual bias was reduced as attention quality increased by attention steering” p22: “Existing issues: insufficient description about reasonability matrix, Ill-formed local explanation visualization…
Suresh et al. (2022) How AI Works: Isolated presentation with interactivity;
Why AI Merits Use: Explanation as retriever
Logical Effects of explanation on human-AI agreement:
p776: We recorded the percent of cases in which participants agreed with the model (versus when they disagreed or were not sure). For cases in which the prediction was correct, the agreement rate was similar across conditions; however, when the prediction was incorrect, we found that participants were less likely to accept the model’s prediction when they were using the NN interface, with or without the input editor. Often in these cases, they did not explicitly “disagree” with the model, but wanted additional information about the signal and/or patient before committing to an answer.
Qualitative feedback:
p776: “Nearest neighbors enable reasoning with clinically-relevant concepts…In contrast, with the baseline condition, participants often had difficulty extracting higher-level, clinically-relevant concepts from the feature importance visualization…Nearest neighbors help characterize uncertainty and incorporate it into decision-making
van der Waa et al. (2021) How AI Works: Isolated presentation; Why AI Merits Use: Explanation as recommender
Logical Identification of model decisive factor and advice:
p15: “First, rule-based explanations indeed seem to allow par- ticipants to more accurately identify the factor from a situation that was decisive in the system’s advice. However, rule-based nor example-based explanation allowed participants to learn to predict system behavior.
Effects of explanations on perceived system understanding and prediction of correctness:
p15: “The rule-based explanations however, did cause to participants to think that they better understood the system compared to example-based and no explanations. The example-based explanations only showed a small and insignificant increase in perceived system understanding.
Vasconcelos et al. (2023) How AI Works: Isolated presentation
Logical Effects of explanations on overreliance and the influence of task difficulty:
p14: “an increase in task difficulty increases the cost of following the strategy to engage with the task, as it generally takes more cognitive effort for people to either (a) ensure that the AI has gotten the answer correct when provided with only predictions or (b) complete the task alone. However, the strategy to rely on the AI does not incur the same costs, as people are not engaging with how hard the task is.” p21: “the salient explanation condition has less overreliance than all other conditions (prediction, highlight explanation, written explanation, and incomplete explanation…We also find that the incomplete explanation condition has less overreliance than the prediction only and written explanation conditions, more overreliance than the salient explanation condition and finds no difference in overreliance in the highlight explanation condition.).”
Wambsganss et al. (2020) How AI Works: Adaptive and Mechanistic presentation;
Why AI Merits Use: Explanation as recommender
Logical Self-reported intention to use, perceived usefulness and ease of use:
p9: “the mean value of intention to use of participants using AL as a writing tool was 2.33 (SD= 0.58). These values are significantly better than the results of the alternative scripting approach. For perceived usefulness we observed a mean value of 3.2 (SD= 1.12) and for perceived ease of use the value was 2.83 (SD= 1.08) for participants from the control group. The mean value for the intention to use was 3.5 (SD= 1.13). The results clearly show that the participants of our experiment rated the acceptance of AL as an adaptive feedback tool positively compared to the usage of the alternative application. The statistical significance was also proven in a double-sided t-test for all three constructs (see table 5).
Qualitative feedback:
p9: “The general attitude for AL was very positive. Especially the fast and direct feedback, the graph-like visualization of the argumentation structure and the summarizing scores were mentioned quite often. However, sometimes AL was not correctly classifying claims and premises, which users suggest to improve.
Wang et al. (2019) How AI Works: Mechanistic presentation;
Why AI Merits Use: Explanation as recommender
Logical Qualitative feedback:
p10-11: “Finding Alternative Hypotheses. We found many instances of users seeking alternative hypotheses with or without various types of XAI…P4 found counterfactual reasoning useful even though one cannot change a patient’s history, ’because it’s a matter of possibility. It’s how you rank your differentials…
Credible Qualitative feedback:
p11: “Lack of Trust: Verify with Supplementary, Situational, Source Data. A lack of trust drove users to want to verify the AI’s decision by looking at raw data, in this case, patient case and physical exam notes, or recorded data…without the context of the patient, [she] can’t diagnose based on the vitals itself. So [she] wouldn’t really trust
Wang and Yin (2021) How AI Works: Isolated presentation
Logical Effects of explanations on human understanding of model behaviors and calibrated trust:
p325-326: “The only positive effect of model explanation that we have consistently observed across different decision making tasks is that feature importance explanations increase people’s objective understanding of an AI model, while feature contribution explanations increase people’s subjective understanding of an AI model…feature contribution explanation seems to be able to satisfy more desiderata of AI explanations when people have some domain expertise in the decision making task…For the two example-based explanations in our study, we found minimal evidence on their capability to support trust calibration.
Wang and Yin (2023) How AI Works: Adaptive presentation
Logical Effects of explanations and model updates on perceived system understanding:
p8: “We find that participants’ perceived change of explanations increased as the the explanations of the updated model in Phase 2 became more dissimilar from those of the old model used in Phase 1. In other words, participants in our experiment could perceive the change in model explanations brought up by a model update.
Effects of explanations and model updates on objective just (frequency of accepting model decisions):
p8: “participants’ objective trust gain and subjective trust gain in the AI model from Phase 1 to Phase 2, both across all participants and within subgroups of participants with different levels of understanding of the AI model. However, we find that neither participants’ objective trust nor their subjective trust seems to be affected by the similarity level of model explanations between Phase 1 and Phase 2.
@@
Credible Effects of explanations and model updates on subjective trust gain and satisfaction gain:
See above excerpt about objective trust; p8: “Figure 3(d) further shows participants’ subjective satisfaction gain from Phase 1 to Phase 2, conditioned on their understanding score. Still, participants did not seem to significantly change their satisfaction with the AI model as the similarity of model explanations before and after the update varied.
Wang et al. (2023b) How AI Works: Mechanistic presentation, Emotional cues; Why AI Merits Use: Explanation as recommender
Logical Effects of counterfactual explanations on several tool useful measures: ‘easy to use,’ ‘easy to understand,’ ‘help understand model,’ ‘help find ways to improve plans,’ and ‘will apply loan again’:
p11: “(A) Participants thought GAM Coach was relatively easy and enjoyable to use, and the tool helped them identify actions to obtain a preferred ML decision. (B) All interaction techniques, especially experimenting with hypothetical values, were rated favorably.
Qualitative feedback:
p9: Three main reasons that participants reported choosing plans were that the plans were (1) controllable, (2) requiring small changes or less compromise, or (3) beneficial for life in general…Participants’ explanations for not choosing a plan mostly complemented the reasons for choosing a plan…Preference configuration is helpful…Participants’ explanations for not choosing a plan mostly complemented the reasons for choosing a plan. Some participants also skipped plans because they were puzzled by counterintuitive suggestions, did not understand the suggestions, or just wanted to see more alternatives.
Emotional Effects of explanations on enjoyable to use (See previous excerpt on tool usefulness in logical)
Yang et al. (2020) How AI Works: Isolated presentation
Logical Effects of explanation and spatial layout on user appropriate trust of AI predictions:
p195: “The results support that all the visual explanations largely increase appropriate trust, reduce overtrust, and help users gain more confidence in their decisions…image-based explanations help correct undertrust, but the effect of rose-based explanations on undertrust is inconclusive…For the image-based explanations, all the three spatial layouts lead to a similar level of users’ trust; appropriate trust, overtrust, and undertrust are very similar and display few differences among the three spatial layouts
Effects of explanation on perceived helpfulness:
Helpfulness was measured but not directly reported in results
@@
Credible Effects of explanation on self-confidence, trust meter before and after AI feedback:
p196-197: “Before seeing the feed- back, participants increase the trust meter for a correct recommendation and decrease the trust meter for an incorrect one; having an explanation shows small positive effects on trust direction…Image-based explanations outperform rose-based explanations because they increase appropriate trust, decrease overtrust and undertrust, improve self-confidence, and show more usability
Zhang et al. (2020) How AI Works: Isolated presentation
Logical Effects of explanation on human-AI agreement:
The baseline condition (green) and the explanation condition (blue) had similar agreement percentages, while the with-confidence condition (orange) had higher percentage when the confidence level was above 70%.